JammarWalker:

More Math HaHa Angles X and Y are supplementary. Angle X is 3 times the measure of angle Y. What is the measure of angle X? 45° 60° 120° 135°

darkknight:

both angles are supplementary, add to 180 degrees so if X = 3Y and we know X + y = 180 we can say 3y+y=180 once we find y we can plug into this equation (x = 3y) to find x

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight both angles are supplementary, add to 180 degrees so if X = 3Y and we know X + y = 180 we can say 3y+y=180 once we find y we can plug into this equation (x = 3y) to find x \(\color{#0cbb34}{\text{End of Quote}}\) oh ok

darkknight:

or we can do x = 3y same thing as x/3 = y so x + x/3 = 180 and solve for x. either way works

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight or we can do x = 3y same thing as x/3 = y so x + x/3 = 180 and solve for x. either way works \(\color{#0cbb34}{\text{End of Quote}}\) uh so its 135?

darkknight:

yes, congrats

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight yes, congrats \(\color{#0cbb34}{\text{End of Quote}}\) Thanks lad.